OpenStudy (anonymous):

sum of 1+3+5+7...... upto n terms

OpenStudy (science0229):

Let's write it like this: 1 +3+ 5 +...+(2n-1)=S (2n-1)+(2n-3)+(2n-5)+...+1=S

OpenStudy (science0229):

notice that the nth term of the sequence is 2n-1

OpenStudy (science0229):

right?

OpenStudy (science0229):

@shreeya13